OpenStudy (anonymous):

Find the area of triangle with given vertices. (1/2) ||u x v||. A(0,0,0), B(1,0,3), C(-3,2,0)

OpenStudy (anonymous):

oh and this involves vectors

OpenStudy (anonymous):

I've already gotten the correct answer, my only question is when you go about using the cross product, can I choose any vector combination AB x AC, BA x BC...etc.?

OpenStudy (amistre64):

the cross of 2 vectors gives us a new vector; and the magnitude of that new vector is the "area" of the parallelpiped

OpenStudy (amistre64):

you should be able to yes

OpenStudy (amistre64):

since A is at 0, your B and C are vector components already

OpenStudy (amistre64):

BxC = say N |N| = area of the parallelagram; so half that is the triangle part

OpenStudy (anonymous):

ok, cause I noticed that no matter which one I choose (AB x AC or BA x BC) the magnitude for both were the same.

OpenStudy (amistre64):

x 1 -3 x = 0-6 y 0 2 -y = (0--9) z 3 0 z = 2-0 <-6,-9, 2> ^2 = \(\sqrt{36+81+4}\)/2 for triangle area

OpenStudy (amistre64):

they are all the same parts of the same triangle, so they should match up in the end

OpenStudy (anonymous):

okay. Thanks.